# Focus on This, Not That! Steering LLMs With Adaptive Feature Specification

### v3

@@ Appendix I Spurious Sentiment (SS) Dataset
We generate a tertiary NLI dataset, SMNLI, with a known spurious feature. We do this considering the MNLI dataset (Williams et al., 2018). This is a NLI dataset with three labels: entailment (0), neutral (1) and contradiction (2), where data is sampled from 5 underlying categories or genres (telephone, government, travel, fiction or slate). We aim to induce spurious correlations between the underlying genres and labels.
@@
Data-generating process (DGP). We consider a graphical model to describe the DGP of examples within the SMNLI dataset. We use the following variables within our DGP:
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original dataset.
@@
  • $S$ spurious feature, here this is the genre of the premise and hypothesis. This is a categorical variable.
  • $X$ example from the MNLI dataset.
  • $Y$ final label for element $X$.
@@
Figure 13: SMNLI DGP. Graphical model showing the data generating process for modifying examples from the MNLI dataset to introduce a new spurious keyword feature $S$.
@@
More specifically, given the orignal dataset $\mathcal{D}$ that we are sub-sampling from, the functions that we use within the DGP for the SMNLI dataset are given by:
$S$ $\sim\texttt{Cat}(\mathcal{S}),$ (42)
$U_{C}$ $\sim\texttt{Ber}(\rho_{\text{spurious}});$ (43)
$C$ =$=\begin{cases}y_{S}&\text{ if }U_{C}=1;\\ \sim\texttt{Unf}(\text{Val}(Y)\setminus\{y_{S}\})&\text{ if }U_{C}=0;\end{cases};$ (44)
@@
Here, $\texttt{Cat}(\mathcal{S})$ is a uniform categorical distribution over spurious feature values (here the underlying genre of the premise-hypothesis pairs). Furthermore, we define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(x|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and genre $S$.
We restrict the genres within the model to $S\in\{\textit{slate},\textit{government},\textit{fiction},\textit{travel}\}$, a subset of the genres of the training set. When creating a distribution shifted test set, we restrict the genres to $S\in\{\textit{facetoface},\textit{nineeleven},\textit{verbatim}\}$. The specific spurious correlations between a genre $s$ and a label $y_{s}$ are chosen to be: $y_{\textit{slate}}=0$; $y_{\textit{government}}=2$; $y_{\textit{fiction}}=1$; $y_{\textit{travel}}=0$; $y_{\textit{facetoface}}=2$; $y_{\textit{nineeleven}}=0$; $y_{\textit{verbatim}}=1$.
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between $S$ and $Y$.
We show that the notion of $\rho_{\text{spurious}}$ in Equation 43 aligns with the notation in Definition 3.1.
Proposition J.1.
From the SCEs described above, we have that
@@
SCM for SMNLI. The DGP again induces a SCM that induces spurious correlations between spurious features $S$ and the label $Y$. The SCM has the same structure as in the SS dataset, and is given in Figure 16 where once again, $U$ again is some hidden confounding variable.
Figure 14: SMNLI SCM. SCM showing showing the spurious correlation present between the keyword feature $S$ and the label $Y$ of examples within the SMNLI dataset, induced through the described sub-sampling process of the MNLI dataset.
Independence conditions during training for FIT. As specified in Appendix C, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features, where, again, $Y_{S}$ is the label spuriously associated to spurious feature value $S$. The results below give sufficient conditions for this to occur with respect to the DGP described in Figure 13.
Proposition J.2.
Assuming the SCEs given above and the corresponding DGP described in Figure 13, if we have a uniform label distribution $p(Y=y)$, and have that $\rho_{\text{spurious}}=1/3$ in the SMNLI training set, then we have that $Y\perp\!\!\!\perp S$.
@@
From Proposition J.1, we have that $\mathbb{P}(Y=y_{s}\mid S=s)=1/3$ when $\rho_{\text{spurious}}=1/3$ for any spurious feature value $s\in\text{Val}(S)$. Moreover, we have that from Equation 44
$\mathbb{P}(Y=y\mid S=s)$ $=\mathbb{P}(C=y\mid S=s)$ (48)
$=\underbrace{\mathbb{P}(U_{C}=1)}_{=\frac{2}{3}}\underbrace{\mathbb{P}(C=Y|S=s)}_{=\frac{1}{2}}$ (49)
$=\frac{1}{3},$ (50)
for each $y\in\text{Val}(Y)\setminus\{y_{s}\}$.
Using this alongside the assumed balanced label distribution, $\mathbb{P}(Y=y)=1/3$ for $y\in\text{Val}(Y)$, implies that we have $\mathbb{P}(Y=y|S=s)=\mathbb{P}(Y=y)$ for all $y\in\text{Val}(Y)$ and $s\in\text{Val}(S)$, which gives that $Y\perp\!\!\!\perp S$. ∎
Proposition J.3.
Assuming the SCEs given above and the corresponding DGP described in Figure 13, if we have a uniform label distribution $p(Y=y)$, and have that $\rho_{\text{spurious}}=1/3$ in the SMNLI training set, then we have that $Y_{S}\perp\!\!\!\perp C$.
@@
First, using that $\rho_{\text{spurious}}=1/3$, we have from Equation 44 that
$\mathbb{P}(C=c\mid S=s)$ =$=\begin{cases}1/3&\text{if }c=f(s),\\ 2/3\cdot 1/2&\text{else.}\end{cases}$ (51)
$=\frac{1}{3},$ (52)
[…]
Now we consider the marginal distribution of C. Let $|\text{Val}(S)|=k$ be the number of spurious feature values. Using the previous result, we have that:
$\mathbb{P}(C=c)$ $=\sum_{s}\underbrace{\mathbb{P}(C=c\mid S=s)}_{=1/3}\underbrace{\mathbb{P}(S=s)}_{1/k}$ (53)
$=\sum_{s}\frac{1}{3k}$ (54)
$=\frac{1}{3}.$ (55)
Therefore, we have that $\mathbb{P}(C=c\mid S=s)=\mathbb{P}(C=c)$, for all $c\in\text{Val}(C)$ and $s\in Val(S)$, which then implies that $S\perp\!\!\!\perp C$. Therefore, using that the spurious label $Y_{S}$ is a deterministic function of $S$, we have that $Y_{S}\perp\!\!\!\perp C$ under the assumptions within the proposition. ∎
Appendix K Spurious HANS dataset (SHANS)
@@
We generate a binary NLI dataset, SHANS, with a known spurious feature. We do this considering the HANS dataset (McCoy, 2019). This is an NLI data set with two labels: entailment (0) and contradiction (1). This is an adversarial dataset designed to assess different NLI models’ reliance on spurious heuristics rather than on the underlying relationship between the premise and the hypothesis when making predictions. Specifically, the author’s consider three major categories of heuristics: lexical overlap heuristic (assuming that a premise entails from words within the hypothesis) , sub-sequence heuristic( assuming that the premise entails all any of its contiguous sub-sequences of words) and constituent heuristic (assuming that a premise entails a hypothesis that is any constituent within it’s syntactic parse tree).
@@
Data-generating process (DGP). We consider a graphical model to describe the DGP of examples within the SHANS dataset. We use the following variables within our DGP:
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original dataset.
@@
  • $Y$ final label for element $X$.
Figure 15: SHANS DGP. Graphical model showing the data generating process for modifying examples from the SHANS dataset to introduce a new spurious features $S_{\text{lex.}}$, $S_{\text{sub.}}$ and $S_{\text{const.}}$ which are encoded within the categorical spurious feature $S$ which represents one of these three heuristics.
@@
where here we define $S$ to be a categorical feature over the set of the presence of each of the three heuristics introduced above which we denote, through overloaded notation, by $\mathcal{S}=\{s_{\text{lex.}},s_{\text{sub.}},s_{\text{const.}}\}$. More specifically, given the original dataset $\mathcal{D}$ that we are sub-sampling from, the functions that we use within the DGP for the S-HANS dataset are given by:
$S$ $\sim\texttt{Cat}(\mathcal{S}),$ (60)
$U_{C}$ $\sim\texttt{Ber}(\rho_{\text{spurious}});$ (61)
$C$$\sim\begin{cases}y_{S}&\text{ if }U_{C}=1;\\ \sim U(\text{Val}(Y)\setminus\{y_{S}\})&\text{ if }U_{C}=0;\end{cases};$ (62)
@@
Here, $\texttt{Cat}(\mathcal{S})$ is a uniform categorical distribution over $\mathcal{S}$ which effectively selects the precence of exactly one of the three spurious feature heuristics. We define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(x|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and the presence of spurious heuristics $S$.
@@
We consider the presence of each feature to be separate binary spurious features. The specific spurious correlations between heuristics and labels $Y$ are chosen to be: $y_{S_{\text{lex.}}=1}=0$; $y_{S_{\text{sub.}}=1}=0$; $y_{S_{\text{const.}}=1}=1$.
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between the heuristics and $Y$.
@@
Transferred results from SMNLI. As we use the same DGP as for the SMNLI dataset described in Appendix K, all of the results that we have proven for SMNLI, translate to the SHANS dataset. In particular, we have that $\rho_{\text{spurious}}$ aligns with the notation in Definition 3.1, and that we have $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ under the assumptions of balanced label distributions and $\rho_{\text{spurious}}=1/2$ within the training set used for FIT training.
SCM for SHANS. The DGP again induces a SCM. In particular, considering $S$ as consiting of three binary spurious features $s_{\text{lex.}}$, $s_{\text{sub.}}$ and $s_{\text{const.}}$. The SCM has a similar structure to as in the SS and S-MNLI datasets, and is given in Figure 16 where once again, $U$ again is some hidden confounding variable.
Figure 16: SHANS SCM. SCM showing the spurious correlations present between the binary precence of heurstics features $S_{\text{lex.}}$, $S_{\text{sub.}}$ and $S_{\text{const.}}$ and the label $Y$ of examples within the S-HANS dataset, induced through the described sub-sampling process of the S-HANS dataset.
Appendix L FIT on SHANS
Here we give the results of performing SFT and FIT on the SHANS datasets.

### v4

@@ Appendix F Spurious NLI dataset (SMNLI)
We generate a tertiary NLI dataset, SMNLI, with a known spurious feature. We do this subsampling from the original MNLI dataset (Williams et al., 2018). This is a NLI dataset with three labels: entailment (0), neutral (1) and contradiction (2), where data is sampled from 6 underlying categories or genres (government, travel, and fiction, facetoface, nineeleven and verbatim). We aim to induce spurious correlations between the underlying genres and labels.
F.1 Data-Generating Process (DGP).
@@
We consider a graphical model to describe the DGP of examples within the SMNLI dataset. We use the following variables within our DGP:
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original MNLI dataset.
  • $X$ example premise and hypothesis from the MNLI dataset.
@@
  • $S$ spurious feature present in the example $X$, here this is the genre of the premise and hypothesis texts. This is a categorical variable.
  • $Y$ final label for example $X$.
@@
Figure 13: SMNLI DGP. Graphical model showing the data generating process for subsampling examples from the MNLI dataset to introduce a new spurious keyword feature $S$.
@@
More specifically, given the original dataset $\mathcal{D}$ that we are sub-sampling from, the functions that we use within the DGP for the SMNLI dataset are given by:
$S$ $\sim\mathcal{U}(\mathcal{S}),$ (38)
$U_{C}$ $\sim\text{Ber}(\rho_{\text{spurious}});$ (39)
$C$ $=U_{C}\;y_{S}+(1-U_{C})\;\mathcal{U}(\text{Val}(C)\setminus\{y_{S}\})$ (40)
@@
Here, $\mathcal{U}(\mathcal{S})$ denotes a uniform distribution over the set of genres $\mathcal{S}$, and we define $\text{Val}(C)=\{0,1,2\}$ corresponding to the possible NLI labels for this task. Furthermore, we define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(\cdot|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and genre $S$.
We restrict the genres that we sample from to $S\in\{\text{government},\text{fiction},\text{travel}\}$, a subset of the genres of the training set. When creating a distribution shifted test set, we restrict the genres to $S\in\{\text{facetoface},\text{nineeleven},\text{verbatim}\}$. The specific spurious relationships between genre values $s$ and their associated spurious labels $y_{s}$ are chosen to be: $y_{\text{slate}}=0$; $y_{\text{government}}=2$; $y_{\text{fiction}}=1$; $y_{\textit{travel}}=0$; $y_{\text{facetoface}}=2$; $y_{\text{nineeleven}}=0$; $y_{\text{verbatim}}=1$. In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between $S$ and $Y$.
We show that the notion of $\rho_{\text{spurious}}$ in Equation 39 aligns with the notation in Definition 2 and that this does not depend on the spurious feature value of $S$.
Proposition F.1.
From the DGP described above, we have that
@@
F.2 Independence Conditions During Training for FIT on SMNLI.
As specified in Appendix D, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features, where, again, $Y_{S}$ is the label spuriously associated to spurious feature value $S$. The results below give sufficient conditions for this to occur with respect to the DGP described in Figure 13. This boils down to setting $\rho_{\text{spurious}}=1/3$ during training.
Proposition F.2.
Assuming the DGP described above and that $\rho_{\text{spurious}}=1/3$, we have that $Y\perp\!\!\!\perp S$.
@@
Note that since $Y=C$ in the DGP above, it suffices to show that $C\perp\!\!\!\perp S$. We have that for a given $S=s$, that $\mathbb{P}(C=y_{s}|S=s)=\rho_{\text{spurious}}=1/3$ from Proposition F.1. Moreover, let $\tilde{C}\sim\mathcal{U}(\text{Val}(C)\setminus\{y_{S}\})$ denote the random variable sampled from a uniform distribution over the remaining possible values of $C$ excluding $y_{s}$. Then for either value of $c\in\text{Val}{(C)}\setminus\{y_{s}\}$ we have that
$\mathbb{P}(C=c|S=s)$ $=\mathbb{P}(U_{C}=0)\mathbb{P}(\tilde{C}=c)$ (44)
$=\frac{2}{3}\cdot\frac{1}{2}$ (45)
$=\frac{1}{3}.$ (46)
Therefore, we have that $C\sim\mathcal{U}(\text{Val}(C))$. In particular, this then gives that $C\perp\!\!\!\perp S$, which in turn implies that $Y\perp\!\!\!\perp S$. ∎
Proposition F.3.
Assuming the DGP described above and that $\rho_{\text{spurious}}=1/3$, then we have that $Y_{S}\perp\!\!\!\perp C$.
@@
F.3 Results.
We present full results comparing FIT on SMNLI against all of the baselines described in Section D.4 in Figure 14, Table 5 for completeness.
Figure 14: Complete Baseline vs FIT Focus Accuracies on SMNLI ($\uparrow$). Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SMNLI dataset. The maximum standard deviations of FIT, $\text{SFT}(y_{\text{focus}})$, $\text{SFT}(y)$ and few-shot methods across models are 6.47%, 7.98%, 1.15%, and 0.500% respectively. fcs = focus, ign = ignore.
Table 5: Complete Baselines vs. FIT Focus Accuracies on SMNLI ($\uparrow$). For each method, we report the mean of focus‐accuracy means ($\mathcal{A}_{\text{focus}}$) over the four test splits ($\mathcal{D}_{\text{iid}}$, $\mathcal{D}_{\text{high}}$, $\mathcal{D}_{\text{low}}$ and $\mathcal{D}_{\text{flipped}}$) $\pm$ the between‐split standard deviation of these means (i.e. how performance varies as predictivity changes) across repeats. Boldface indicates the best mean for each focus instruction type and model independently.
$\varnothing$ $\mathrm{focus}(C)$ $\mathrm{focus}(C)\wedge\mathrm{ignore}(S)$ $\mathrm{ignore}(S)$ $\mathrm{focus}(S)$ $\mathrm{focus}(S)\wedge\mathrm{ignore}(C)$
Mistral Zero-shot $0.742_{\pm 0.016}$ $0.711_{\pm 0.014}$ $0.711_{\pm 0.012}$ $0.716_{\pm 0.020}$ $0.362_{\pm 0.189}$ $0.368_{\pm 0.191}$
Few-shot $0.716_{\pm 0.020}$ $0.725_{\pm 0.017}$ $0.722_{\pm 0.017}$ $0.687_{\pm 0.023}$ $0.370_{\pm 0.178}$ $0.386_{\pm 0.189}$
SFT($y$) $\mathbf{0.890_{\pm 0.009}}$ $0.875_{\pm 0.013}$ $\mathbf{0.875_{\pm 0.013}}$ $\mathbf{0.885_{\pm 0.013}}$ $0.334_{\pm 0.275}$ $0.332_{\pm 0.273}$
SFT($y_{\mathrm{focus}}$) $0.732_{\pm 0.101}$ $0.727_{\pm 0.097}$ $0.725_{\pm 0.096}$ $0.726_{\pm 0.102}$ $0.544_{\pm 0.192}$ $0.537_{\pm 0.190}$
FIT $0.878_{\pm 0.005}$ $\mathbf{0.875_{\pm 0.004}}$ $0.874_{\pm 0.006}$ $0.878_{\pm 0.007}$ $\mathbf{0.963_{\pm 0.006}}$ $\mathbf{0.968_{\pm 0.003}}$
Llama Zero-shot $0.679_{\pm 0.007}$ $0.688_{\pm 0.017}$ $0.682_{\pm 0.018}$ $0.682_{\pm 0.015}$ $0.370_{\pm 0.162}$ $0.386_{\pm 0.151}$
Few-shot $0.595_{\pm 0.012}$ $0.641_{\pm 0.003}$ $0.630_{\pm 0.010}$ $0.636_{\pm 0.009}$ $0.351_{\pm 0.130}$ $0.379_{\pm 0.094}$
SFT($y$) $\mathbf{0.880_{\pm 0.005}}$ $\mathbf{0.880_{\pm 0.003}}$ $\mathbf{0.879_{\pm 0.004}}$ $\mathbf{0.879_{\pm 0.004}}$ $0.352_{\pm 0.278}$ $0.354_{\pm 0.272}$
SFT($y_{\mathrm{focus}}$) $0.700_{\pm 0.121}$ $0.759_{\pm 0.093}$ $0.750_{\pm 0.098}$ $0.720_{\pm 0.110}$ $0.552_{\pm 0.194}$ $0.536_{\pm 0.209}$
FIT $0.872_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.861_{\pm 0.009}$ $0.865_{\pm 0.010}$ $\mathbf{0.929_{\pm 0.005}}$ $\mathbf{0.931_{\pm 0.004}}$
Vicuna Zero-shot $0.595_{\pm 0.012}$ $0.598_{\pm 0.015}$ $0.575_{\pm 0.019}$ $0.573_{\pm 0.017}$ $0.334_{\pm 0.108}$ $0.341_{\pm 0.093}$
Few-shot $0.269_{\pm 0.032}$ $0.417_{\pm 0.027}$ $0.251_{\pm 0.019}$ $0.301_{\pm 0.013}$ $0.195_{\pm 0.080}$ $0.133_{\pm 0.060}$
SFT($y$) $\mathbf{0.892_{\pm 0.003}}$ $\mathbf{0.890_{\pm 0.002}}$ $\mathbf{0.890_{\pm 0.004}}$ $\mathbf{0.891_{\pm 0.006}}$ $0.337_{\pm 0.288}$ $0.342_{\pm 0.283}$
SFT($y_{\mathrm{focus}}$) $0.579_{\pm 0.081}$ $0.575_{\pm 0.064}$ $0.573_{\pm 0.055}$ $0.585_{\pm 0.063}$ $0.502_{\pm 0.092}$ $0.496_{\pm 0.086}$
FIT $0.803_{\pm 0.004}$ $0.807_{\pm 0.006}$ $0.794_{\pm 0.008}$ $0.781_{\pm 0.012}$ $\mathbf{0.894_{\pm 0.019}}$ $\mathbf{0.934_{\pm 0.010}}$
Appendix G Additional BBQ Results
In Figure 15 and Table 6 we include the full comparisons of FIT against all baselines described in Section D.4 on BBQ.
Figure 15: Complete Baseline vs FIT Focus Accuracies on BBQ ($\uparrow$). Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the BBQ dataset. The maximum standard deviations of FIT, $\text{SFT}(y_{\text{focus}})$, $\text{SFT}(y)$ and few-shot methods across models are 4.07%, 10.7%, 2.22%, and 0.600% respectively. fcs = focus, ign = ignore.
Table 6: Complete Baseline vs FIT Focus Accuracies on BBQ ($\uparrow$). We report across all baselines the mean seen/unseen focus accuracies, with boldface indicating the best mean for each focus instruction type and model independently. The maximum standard deviations of FIT, $\text{SFT}(y_{\text{focus}})$, $\text{SFT}(y)$ and few-shot methods across models are 4.07%, 10.7%, 2.22%, and 0.600% respectively. fcs = focus, ign = ignore.
$\varnothing$ $\mathrm{focus}(C)$ $\mathrm{focus}(C)\wedge\mathrm{ignore}(S)$ $\mathrm{ignore}(S)$ $\mathrm{focus}(S)$ $\mathrm{focus}(S)\wedge\mathrm{ignore}(C)$
Mistral Zero-shot 0.772/0.792 0.768/0.803 0.782/0.812 0.790/0.829 0.338/0.320 0.302/0.295
Few-shot 0.775/0.798 0.767/0.795 0.769/0.807 0.777/0.814 0.408/0.381 0.388/0.373
SFT($y$) 0.999/0.982 1.000/0.983 1.000/0.982 1.000/0.981 0.248/0.249 0.248/0.249
SFT($y_{\mathrm{focus}}$) 0.858/0.874 0.844/0.866 0.848/0.872 0.847/0.863 0.396/0.335 0.394/0.340
FIT 0.998/0.976 0.998/0.976 0.998/0.977 0.999/0.976 0.941/0.852 0.941/0.853
Llama Zero-shot 0.534/0.673 0.478/0.597 0.478/0.593 0.508/0.627 0.533/0.421 0.527/0.428
Few-shot 0.643/0.696 0.686/0.743 0.724/0.774 0.642/0.720 0.496/0.429 0.492/0.432
SFT($y$) 0.998/0.991 0.999/0.985 1.000/0.988 0.999/0.989 0.248/0.244 0.248/0.246
SFT($y_{\mathrm{focus}}$) 0.836/0.912 0.825/0.901 0.836/0.903 0.852/0.910 0.403/0.317 0.418/0.323
FIT 0.993/0.974 0.997/0.977 0.998/0.979 0.996/0.979 0.943/0.832 0.946/0.835
Vicuna Zero-shot 0.444/0.495 0.495/0.567 0.505/0.531 0.479/0.519 0.371/0.324 0.375/0.315
Few-shot 0.454/0.565 0.453/0.572 0.517/0.604 0.547/0.625 0.503/0.512 0.487/0.496
SFT($y$) 0.975/0.991 0.959/0.988 0.957/0.990 0.974/0.990 0.245/0.252 0.243/0.251
SFT($y_{\mathrm{focus}}$) 0.796/0.832 0.794/0.808 0.787/0.806 0.796/0.809 0.444/0.392 0.456/0.394
FIT 0.983/0.979 0.982/0.975 0.981/0.972 0.984/0.973 0.966/0.835 0.966/0.837
Appendix H BBQ-NLG
H.1 BBQ-NLG Experimental Setup
We follow the original BBQ dataset splits (including both training and held-out social-bias partitions) described in Section 4.2. For each example in the new BBQ-NLG dataset, we remove the predefined answer choices and require the model to generate fully verbalised responses, rather than selecting from a fixed set of three options as in the original BBQ dataset and outputting a categorical label.
To give the model additional capacity for this more challenging free-form generation task, we augment the usual LoRA targets (the key and value projection matrices) with adapters on the value-projection matrices as well. All LoRA modules are trained with a learning rate of $10^{-5}$. We train both the FIT and $\text{SFT}(y_{\text{focus}})$ models for 10 epochs each. As a strong comparison, we evaluate a few-shot baseline that conditions the model on 4 in-context examples that share the same focus instruction as the test instance.
To enable both rapid and cost-effective evaluation of correctness in our open-ended generation task, we employ an LLM-based judge. Specifically, we use a pre-trained Llama-3.1-8B-Instruct model to compare each generated response against its associated ground-truth focus label and determine whether they are semantically equivalent. Manual checks confirm that the Llama-3.1-8B-Instruct judge reliably assesses semantic equivalence in this setting where model generations and expected responses and generally short and concise.
H.2 Results
Figure 16 reports the focus-accuracy of each method across models. Although all approaches exhibit a slight degradation on unseen feature combinations relative to the classification-style BBQ results in Section 4.2, FIT remains generally overall on par with the earlier numbers. Critically, FIT consistently outperforms both the $\text{SFT}(y_{\text{focus}})$ and the strong 4-shot few-shot baseline that uses identical focus prompts at test time. These findings provide clear evidence that the FIT method can be effectively extended to open-ended NLG settings without sacrificing its steering capabilities.
Figure 16: BBQ-NLG Focus Accuracies ($\uparrow$). Mean focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the BBQ-NLG dataset. The maximum standard deviations of FIT, $\text{SFT}(y_{\text{focus}})$ and few-shot methods across models and $\mathcal{I}_{\text{focus}}$ are 2.45%, 6.35% and 0.377% respectively. fcs = focus, ign = ignore.
Appendix I Comparison of FIT Against a Specific Debiasing Technique
FIT is a general framework designed to enable users to steer a model’s behaviour based on specified features. This approach provides enhanced control over model outputs during inference, adding a critical layer of explainability and controllability to model predictions.
Results. The results of the debiasing experiment comparing FIT to the PoE method is shown in Figure 17. FIT performs equally as well as the PoE method as shown by comparing the default prompt accuracy ($\emptyset$) for the PoE models against the $\text{focus}(C)$ results for the FIT models; both metrics correspond to causal accuracy for these prompt types. This indicates that FIT performs just as well as the PoE dedicated debiasing technique.
However, the PoE method requires training two separate models and does not provide steerability at test time as shown by the low focus accuracy on $\text{focus}(S)$. Indeed the model defaults to the ground truth label across all prompt types and does not change behaviour in the presence of different focus specifications. This highlights the flexibility of FIT, which not only debiases effectively but also enables additional controllability during inference.
Figure 17: Focus Accuracy of FIT Against PoE Debiasing Technique ($\uparrow$). Figure showing the focus accuracies ($\mathcal{A}_{\text{focus}}$) of FIT (bottom row) and the dedicated debiasing technique, PoE (top row), on the BBQ dataset.
Appendix J Spurious HANS Dataset (SHANS)
@@ Appendix I Comparison of FIT Against a Specific Debiasing Technique
We generate a binary NLI dataset, SHANS, with a known spurious feature. We do this by subsampling from the HANS dataset (McCoy et al., 2019). This is an NLI data set with two labels: entailment (0) and contradiction (1). This is an adversarial dataset designed to assess different NLI models’ reliance on spurious heuristics rather than on the underlying relationship between the premise and the hypothesis when making NLI predictions. Specifically, the author’s consider three major categories of heuristics: lexical overlap heuristic (assuming that a premise entails from words within the hypothesis) , sub-sequence heuristic (assuming that the premise entails all any of its contiguous sub-sequences of words) and constituent heuristic (assuming that a premise entails a hypothesis that is any constituent within it’s syntactic parse tree). We take each of these as separate spurious features within our SHANS dataset for which we induce spurious correlations, as for the SMNLI dataset, through subsampling.
J.1 Data-Generating Process (DGP).
@@
We consider a graphical model to describe the DGP of examples within the SHANS dataset. We use the following variables within our DGP:
@@
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original HANS dataset.
@@
  • $Y$ final label for example $X$.
Figure 18: SHANS DGP. Graphical model showing the DGP for subsampling examples from the SHANS dataset to introduce new spurious features $S_{\text{lex.}}$, $S_{\text{sub.}}$ and $S_{\text{const.}}$ which are encoded within the categorical spurious feature $S$ which represents one of these three heuristics.
@@
where here we define $S$ to be a categorical feature over the set of variables indicating the presence of each of the three heuristics introduced above which we denote, through overloaded notation, by $\mathcal{S}=\{s_{\text{lex.}},s_{\text{sub.}},s_{\text{const.}}\}$. More specifically, given the original HANS dataset $\mathcal{D}$ that we are sub-sampling from, the functions that we use within the DGP for the SHANS dataset are given by:
$S$ $\sim\mathcal{U}(\mathcal{S}),$ (51)
$U_{C}$ $\sim\text{Ber}(\rho_{\text{spurious}});$ (52)
$C$ $\sim U_{C}y_{S}+(1-U_{C})(1-y_{S});$ (53)
@@
Here, $\mathcal{U}(\mathcal{S})$ is a uniform categorical distribution over the set of spurious features $\mathcal{S}$ which effectively selects the presence of exactly one of the three spurious feature heuristics. We define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(\cdot|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and the presence of spurious heuristic $S$.
We restrict the genres to $S\in\{s_{\text{lex.}},s_{\text{const.}}\}$ for heuristics seen during training. We then create additional test sets with spurious feature set restricted to $S\in\{s_{\text{sub.}}\}$, which serves as an unseen spurious feature set to test generalisation.
@@
We consider the presence of each feature to be separate binary spurious features. The specific spurious correlations between heuristics and labels $Y$ are chosen to be: $y_{S_{\text{lex.}}=1}=0$; $y_{S_{\text{sub.}}=1}=0$; $y_{S_{\text{const.}}=1}=1$. In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between the heuristics $S$ and label $Y$.
J.2 Transferred Results from SMNLI.
@@
As we have effectively used the same DGP as for the SMNLI dataset described in Appendix F, with the only change being the label set, all of the results that we have proven for SMNLI, translate to the SHANS dataset. In particular, we have that $\rho_{\text{spurious}}$ aligns with the notation in Definition 2, and that we have $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ under the assumptions of $\rho_{\text{spurious}}=1/2$ within the training set.
Appendix K FIT on SHANS
Here, we give the results performing FIT on the SHANS dataset.
K.1 Spurious HANS (SHANS) Dataset.